Answer the following questions based on the relationship between real and nominal variables. Round answers to two decimal places as needed. When Joe started his job at the laundromat 5 years ago, his wage was $5.50 an Number hour. Today his wage is $7.00 an hour. If Joe started his job in the base year, and tos rea weis the same as when he started, what is thé Consumer Price lIndexD today? The International Disc Jockeys Union has a wage contract that stipulates a yearly wage increase based on the Consumer Price Index. If this years wage is $27.00, the current CPl is 160, and theontract was first signed in the base year, what was the original salary the first year of the contract? Number $ Shelby bought her dream car, a 1966 red convertible Mustang, with a loan from her Number credit union. If Shelby paid 4.9% and the bank earned a real rate of return of 3.7%, what was the inflation rate over the life of the loan? Number Tim is feeling magnanimous, and decides to loan his mother enough for her to pay off her creditcard bills. If Tim wants to earn 2.5% and expects inflation to be 4.9%, how much will he charge his mother for the loan?
Answer #1

1) CPI = Wage (today)/ Wage (base year) x 100 = 7/5.5 x 100 = 127.27

2) Original salary (base year) = 27 x 100/160 = $16.88

3) Nominal rate of return = Real rate of return + inflation rate

inflation rate = 4.9% - 3.7% = 1.2%

4) Nominal rate of return (interest rate) = 4.9% + 2.5% = 7.4%
Answer #2

olution uestion - Consume Price Inden lCP 6.6D Base pesi od tos CP 127- 27 Question 2 Consumer pnce Irdex Current yeas SaryQuestlon 3 Infla bionÞate2 Nominal Tate 0foe tuzn -Real rate of eburn . 1、2 % Nominal rate of return